OpenStudy (anonymous):

what do 56 and 72 have in common factors

OpenStudy (anonymous):

9 i think

OpenStudy (anonymous):

its 8

OpenStudy (anonymous):

56=7*8 72=8*9

OpenStudy (anonymous):

yeh, that was just a guess with no real thought lol I have seen those numbers so many times before , so I just kinda guessed it lols